OpenStudy (anonymous):

The volume of a sphere is 5,000pi m3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (kinggeorge):

The volume of a sphere is given by \[V={4 \over 3}\pi r^3\]and the surface area is given by \[A=4 \pi r^2\]Since we have volume, we can backtrack to find the radius, and then the surface area. So, \[5000\pi={4 \over 3}\pi r^3\]\[3750=r^3\]\[r=\sqrt[3]{3750}\]Now, we can substitute in to find the surface area.\[A=4\pi (\sqrt[3]{3750})^2\]We can evaluate this with a calculator to get 3033 m^2

OpenStudy (kinggeorge):

To be more precise, we get \[V\approx 3033.1743577959768605787928419 \;\;\;m^2\]which rounds to \[V\approx3033\;\;\;m^2\]